The two plates of a capacitor hold +2500 μC and -2500 μC of charge, respectively, when the potential difference is 940 V

What is the capacitance?
Answer #1

Solution)

Given,

Charge, Q=2500*10^-6 C

Voltage, V=940 V

We know,

Capacitance, C=Q/V

Substitute values,

C=(2500*10^-6)/(940)=2.65*10^-6 F (Ans)

==========